OpenStudy (anonymous):

what is the sum of the first 4 terms of the arithmetic sequence in which the 6th term is 8 and the 10th term is 13

Miracrown (miracrown):

so we need to first write what an arithmetic sequence looks like in general general then we can solve for the terms a_n = a_n+1 + (n-1)d this is the general equation we know that a_6 = 8 and a_10 = 13 so a_7, a_8, a_9 all come in between we need to add something to 8 4 times to get to 13 (13-8)/4 = 5/4 = 1.25 a_6 = 8 a_7 = 9.25 a_8 = 10.5 a_9 = 11.75 a_10 = 13 so that works by adding 1.25 to the a_6 term 4 times, we get to 13 so the "d" term is 1.25 we need to back up to get a_1, a_2, a_3 and a_4 might be easier to just subtract d a_5 = 8 - 1.25 = 6.75 a_4 = 6.75 - 1.25 = 5.5 a_3 = 5.5 - 1.25 = 4.25 a_2 = 4.25 - 1.25 = 3 so the sum is: 1.75 + 3 + 4.25 + 5.5 =)

Miracrown (miracrown):

= 6 + 3 + 5.5 = 9 + 5.5 = 14.5

Miracrown (miracrown):

we could solve for a_n-1 a_n-1 = a_n - (n-1)d but again it just backs us up one step at a time

OpenStudy (anonymous):

Well the 6th term is a +5d = 8 10th term a+9d = 13 5 = 4d, d = 5/4 a = 8 -25/4 = 7/4 So your nth term = 7/4+(5/4)(n-1) = 1/4(5n+2) 4th term = (1/4)(20+2) = 22/4, get the the sum = 29/2 = 14.5